How many gallons of pure antifreeze must be mixed with 30 gallons of 15% antifreeze to get a mixture that is 40% antifreeze ?

Answer by stanbon(75887)

How many gallons of pure antifreeze must be mixed with 30 gallons of 15% antifreeze to get a mixture that is 40% antifreeze ?
-----
Equation:
anti + anti = anti
1*x + 0.15*30 = 0.40(x+30)
-----
100x + 15*30 = 40x + 40*30
60x = 25*30
x = 12.5 gallons (amt. of pure anti to add)
==================================
Cheers,
Stan H.
================